# Removal and Selection: Improving RGB-Infrared Object Detection via Coarse-to-Fine Fusion

### v3

@@ 3.4 Removal and Selection Detector (RSDet)
where $\operatorname{MI}$ represents the operation of mutual information. We use cross-entropy ($\operatorname{CE}$) and KL-divergence ($\operatorname{KL}$) to calculate the mutual information Sanghi (2020); Tschannen et al. (2019) between different features in the latent space.
@@
For detection loss, we also use the detection loss $\mathcal{L}_{\text{rpn}}$, $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{cls}}$ same as the Faster R-CNN Ren et al. (2017) to supervise the whole RSDet network training. Thus, the total loss is the sum of these individual losses:
@@
4 Experiments
4.1 Datasets
@@
KAIST Dataset. KAIST Hwang et al. (2015) is a public aligned multispectral pedestrian detection dataset. Due to the problematic annotations in the original dataset, our framework is trained on 8963 image pairs with new annotations Zhang et al. (2019b) and evaluated on 2252 image pairs with the improved annotations Liu et al. (2016a). We perform comparison experiments under ‘All-dataset’ settings Hwang et al. (2015) and evaluate six subsets in terms of pedestrian distance and occlusion level.
@@
FLIR Dataset. FLIR is a paired visible and infrared object detection dataset including daytime and night scenes. Since the images are misaligned in the original dataset, we use the FLIR-aligned dataset Zhang et al. (2020) in our experiment. It has 5142 aligned RGB-IR image pairs, of which 4129 are used for training and 1013 for testing.
@@
LLVIP Dataset. LLVIP Jia et al. (2021) is a strictly aligned visible and infrared object detection dataset for low-light vision. It is collected in low-light environments most of which were captured in very dark scenes. It contains 15488 aligned RGB-IR image pairs, of which 12025 images are used for training and 3463 images for testing.
4.2 Evaluation Metrics
@@
log-average Miss Rate ($\text{MR}^{\text{-2}}$). For KAIST dataset, we employ the standard KAIST evaluation Hwang et al. (2015): log-average Miss Rate (MR) over False Positive Per Image (FPPI) range of $[10^{-2},10^{0}]$ (also denoted as $\text{MR}^{\text{-2}}$). The lower score indicates better performance.
@@
mean Average Precision (mAP). For FLIR and LLVIP datasets, we employ the commonly used object detection metric $\text{mAP}_{50}$ and mAP. The $\text{mAP}_{50}$ metric represents the mean AP under IoU 0.50 and The mAP metric represents the mean AP under IoU ranges from 0.50 to 0.95 with a stride of 0.05. The higher value of two metrics means better performance.
@@
Table 2: Camparison the performance (mAP, in%) on the FLIR and LLVIP dataset. The best results are highlighted in red and the second-place are highlighted in blue.
@@
SSD Liu et al. (2016b) VGG16 65.5 29.6 90.2 53.5
RetinaNet Lin et al. (2017) ResNet50 66.1 31.5 94.8 55.1
Cascade R-CNN Cai and Vasconcelos (2019) ResNet50 71.0 34.7 95.0 56.8
Faster R-CNN Ren et al. (2017) ResNet50 74.4 37.6 94.6 54.5
DDQ-DETR Zhang et al. (2023) ResNet50 73.9 37.1 93.9 58.6 IR
SSD Liu et al. (2016b) VGG16 52.2 21.8 82.6 39.8
RetinaNet Lin et al. (2017) ResNet50 51.2 21.9 88.0 42.8
Cascade R-CNN Cai and Vasconcelos (2019) ResNet50 56.0 24.7 88.3 47.0
Faster R-CNN Ren et al. (2017) ResNet50 64.9 28.9 87.0 45.1
DDQ-DETR Zhang et al. (2023) ResNet50 64.9 30.9 86.1 46.7 RGB
Halfway fusion Liu et al. (2016a) VGG16 71.5 35.8 91.4 55.1
GAFF Zhang et al. (2021) ResNet18 74.6 37.4 94.0 55.8
ProbEn Chen et al. (2022) ResNet50 75.5 37.9 93.4 51.5
LGADet Zuo et al. (2023) ResNet50 74.5 - - -
CSAA Cao et al. (2023b) ResNet50 79.2 41.3 94.3 59.2
@@
4.3 Implementation Details
All the experiments are implemented in the mmdetection Chen et al. (2019) toolbox and conducted on a single GPU NVIDIA GeForce RTX 3090. We use ResNet-50 as the backbone. During training, the batch size is set to 4 and the SGD optimizer is employed with a momentum of 0.9 and the weight decay of $1\times 10^{-4}$. The input image resolution is $640\times 512$ for FLIR and KAIST and $1280\times 1024$ for LLVIP. All experiments are trained for 12 epochs with an initial learning rate of $1\times 10^{-2}$. We use random flipping with a probability of 0.5 to increase input diversity.
Figure 4: The FPPI-MR curves comparisons with the state-of-the-art methods on the KAIST dataset under the ‘All-dataset’ settings Hwang et al. (2015).
4.4 Comparison with State-of-the-Art Methods
KAIST Dataset. The experiments are conducted on the KAIST under ‘All-dataset’ settings Hwang et al. (2015); Li et al. (2018). We compare the proposed RSDet with twelve SOTA multispectral object detection methods on the KAIST dataset. As shown in Table 1, the RSDet performs best on the ‘All’, ‘Day’, and ‘Night’ conditions and the other four of six subsets (‘Near’, ‘Medium’, ‘Far’, ‘None’). And rank second in two subsets (Partial and Heavy). Specifically, the RSDet takes the lead in the subsets of different scales (‘Near’, ‘Medium’, ‘Far’). Especially, in the ‘Far’ subsets, the RSDet surpasses the second-best by about 10%, which is attributed to the mixture of scale-aware experts in the DFS module. Besides, on the whole test dataset (‘All’), the RSDet surpasses the second-best by about 3%, which indicates the satisfactory accuracy of RSDet. The log-average Miss Rate ($\text{MR}^{\text{-2}}$) over the False Positive Per Image (FPPI) curve of all detectors is also shown in Figure 4. The above results prove the superiority of our method.
FLIR and LLVIP Dataset. The quantitative results of the different methods on the FLIR and LLVIP datasets are shown in Table 2. We compare the proposed RSDet with five SOTA multispectral object detection methods and five single-modality object detection methods. From Table 2, it can be observed that the experimental results of fusion methods are generally better than those of single-modal methods in the FLIR dataset. However, due to the limited light conditions, the RGB features interfere with the IR features in the LLVIP dataset, resulting in the detection results of the multispectral detector being inferior to the IR modality. However, our method effectively addresses this issue with the well-designed coarse-to-fine fusion strategy and achieves state-of-the-art performance on the FLIR and LLVIP datasets.
Table 3: Ablation study on each module result (mAP, in%) on the FLIR and LLVIP dataset. The best results are highlighted in bold.
@@ 4.4 Comparison with State-of-the-Art Methods
4.5 Ablation Study
Ablation on each component. To verify the effectiveness of the RSR and DFS modules, we ablate each module from RSDet. As shown in Table 3, after adding the RSR module the performance increases 1.2% mAP. on the FLIR dataset and 1.6% mAP on the LLVIP dataset respectively. Further adding the DFS module, the metrics show a significant increase on the FLIR dataset, the $\text{mAP}_{50}$ increase 4.8%, and the mAP increase 3.0%. On the LLVIP dataset, the experiment result also exhibits similar trends, especially in the mAP metric. The above results indicate the effectiveness of these modules.
Ablation on different K.  We further study the impact of different values of $K$ in the RSR module. We conduct the experiments by modifying the value of $K$ ranging from 100 to 400 on the FLIR and LLVIP datasets. As shown in Figure 6, our detector achieves the highest detection performance when $K=320$. If the value of $K$ is increased or decreased, the performance of the detector will decrease, which also proves the negative impact of the redundant spectrum for the detector.
4.6 Visualization and Discussion
RSR module. We conduct tSNE visualizations Van der Maaten and Hinton (2008) of shared $C_{\text{sha}}$ and specific features ($C_{\text{I-spe}}$ and $C_{\text{V-spe}}$) on FLIR dataset. As shown in Figure 7, we have observed that without the RSR module, there is still a considerable mix of feature points in the shared and specific features, leading to difficulty in selecting the desired specific features in DFS. However, after adding the RSR module, we have noted a significant decrease in the number of mixed features, which verifies that removing redundant spectra is beneficial to feature disentanglement and, thus more effective in obtaining fused features in the DFS module.
Figure 7: tSNE visualization Van der Maaten and Hinton (2008) of the modality-specific and shared features. (a) ‘w/o RSR’ represents without RSR module. (b) ‘w RSR’ represents without RSR module.
@@ 4.6 Visualization and Discussion
5 Conclusion
In this paper, we presented a new coarse-to-fine perspective to fuse visible and infrared modality features. Specifically, a Redundant Spectrum Removal (RSR) module is first designed to coarsely filter out the irrelevant spectrum, and then a Dynamic Feature Selection (DFS) module is proposed to finely select the desired features for the final fusion process. we constructed a new object detector called Removal and Selection Detector (RSDet) to evaluate its effectiveness and versatility. Extensive experiments on three public RGB-IR detection datasets demonstrated that our method can effectively facilitate complementary fusion and achieve state-of-the-art performance. We believe that our method can be applied to various studies in the RGB-IR feature fusion tasks.

### v4

@@ III-D Removal and Selection Detector (RSDet)
where $\operatorname{MI}$ represents the operation of mutual information. We use cross-entropy ($\operatorname{CE}$) and KL-divergence ($\operatorname{KL}$) to approximate equivalent optimize the mutual information [48, 49] between different features in the latent space.
@@
For detection loss, we also use the $\mathcal{L}_{\text{rpn}}$, $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{cls}}$ same as the Faster R-CNN [47] to supervise the training of whole Removal and Selection Detector. Thus, the total loss is the sum of these individual losses:
@@
IV Experiments
TABLE I: Detection results ($\text{MR}^{\text{-2}}$, in%) under ‘All-dataset’ settings of different pedestrian distances, occlusion levels, and light conditions (Day and Night) on the KAIST dataset. The pedestrian distances consist of ‘Near’ (115 $\leq$ height), ‘Medium’ (45 $\leq$ height $<$ 115) and ‘Far’ (1 $\leq$ height $<$ 45), while occlusion levels consist of ‘None’ (never occluded), ‘Partial’ (occluded to some extent up to half) and ‘Heavy’ (mostly occluded). IoU $=0.5$ / $0.7$ is set for evaluation. The best results are highlighted in red and the second-place are highlighted in blue.
$\text{MR}^{\text{-2}}$ (IoU=0.5)
Methods Backbone Near Medium Far None Partial Heavy Day Night All
ACF [25] VGG16 28.74 53.67 88.20 62.94 81.40 88.08 64.31 75.06 67.74
Halfway Fusion [17] VGG16 8.13 30.34 75.70 43.13 65.21 74.36 47.58 52.35 49.18
FusionRPN+BF [50] VGG16 0.04 30.87 88.86 47.45 56.10 72.20 52.33 51.09 51.70
IAF R-CNN [29] VGG16 0.96 25.54 77.84 40.17 48.40 69.76 42.46 47.70 44.23
IATDNN+IASS [28] VGG16 0.04 28.55 83.42 45.43 46.25 64.57 49.02 49.37 48.96
CIAN [51] VGG16 3.71 19.04 55.82 30.31 41.57 62.48 36.02 32.38 35.53
MSDS-R-CNN [18] VGG16 1.29 16.19 63.73 29.86 38.71 63.37 32.06 38.83 34.15
AR-CNN [52] VGG16 0.00 16.08 69.00 31.40 38.63 55.73 34.36 36.12 34.95
MBNet [20] ResNet50 0.00 16.07 55.99 27.74 35.43 59.14 32.37 30.95 31.87
TSFADet [14] ResNet50 0.00 15.99 50.71 25.63 37.29 65.67 31.76 27.44 30.74
CMPD [11] ResNet50 0.00 12.99 51.22 24.04 33.88 59.37 28.30 30.56 28.98
CAGTDet [31] ResNet50 0.00 14.00 49.40 24.48 33.20 59.35 28.79 27.73 28.96
((a))
$\text{MR}^{\text{-2}}$ (IoU=0.7)
Methods Backbone Near Medium Far None Partial Heavy Day Night All
ACF [25] VGG16 79.25 82.96 97.86 87.59 94.61 97.86 88.48 92.47 89.54
Halfway Fusion [17] VGG16 49.59 74.87 97.00 80.35 90.42 94.15 81.31 86.34 83.15
FusionRPN+BF [50] VGG16 35.78 68.82 99.38 76.29 86.80 92.47 76.98 83.71 79.30
IAF R-CNN [29] VGG16 33.75 70.24 98.12 76.74 84.58 93.69 77.02 84.38 79.59
IATDNN+IASS [28] VGG16 45.40 70.85 99.00 78.25 84.51 93.13 80.46 82.32 80.91
CIAN [51] VGG16 38.31 63.98 87.12 70.39 80.95 91.68 72.44 78.92 74.45
MSDS-R-CNN [18] VGG16 35.49 57.95 93.15 68.41 76.23 90.37 69.85 78.52 71.93
AR-CNN [52] VGG16 25.19 53.88 91.72 64.91 73.18 88.70 64.45 77.29 68.64
MBNet [20] ResNet50 16.98 51.21 85.33 60.84 69.59 86.22 63.50 67.76 65.14
TSFADet [14] ResNet50 19.50 49.32 81.90 58.93 72.09 87.10 61.78 68.38 63.85
CMPD [11] ResNet50 19.31 49.69 83.93 59.79 66.64 84.79 61.77 68.83 63.93
CAGTDet [31] ResNet50 20.80 47.40 78.31 56.95 67.39 85.11 60.24 65.45 61.71
RSDet (Ours) ResNet50 12.68 46.73 73.95 55.35 65.03 84.08 59.72 62.46 60.65
((b))
IV-A Datasets
Figure 5: The FPPI-MR curves comparisons with the state-of-the-art methods on the KAIST dataset under the ‘All-dataset’ settings [25].
IV-A1 KAIST Dataset
@@
KAIST [25] is a public multispectral pedestrian detection dataset. Due to the problematic annotations in the original dataset, our framework is trained on 8963 image pairs with new annotations [52] and evaluated on 2252 image pairs with the improved annotations [17]. To comprehensively evaluate the performance of our method, we perform comparison experiments under ‘All-dataset’ settings [25] and six subset datasets in terms of different pedestrian distances containing “Near”, “Medium”, and “Far” and different occlusion levels containing “None”, “Partial”, and “Heavy”.
IV-A2 FLIR-aligned Dataset
@@
FLIR is a paired visible and infrared object detection dataset including daytime and night scenes. Since the images are misaligned in the original dataset, we use the FLIR-aligned dataset [26]. It has 5142 aligned RGB-IR image pairs, of which 4129 are used for training and 1013 for testing, and contains three classes of objects: ’person’, ’car’, and ’bicycle’. Since there are very few instances of the ’dog’ category in the FLIR-aligned dataset, we clean the annotations and remove the ’dog’ category from the dataset.
IV-A3 LLVIP Dataset
@@
LLVIP [27] is a strictly aligned visible and infrared object detection dataset for low-light vision. It is collected in low-light environments most of which were captured in very dark scenes. It contains 15488 aligned RGB-IR image pairs, of which 12025 images are used for training and 3463 images for testing.
IV-B Evaluation Metrics
IV-B1 Log-average Miss Rate ($\text{MR}^{\text{-2}}$)
@@
For the KAIST dataset, we employ the standard KAIST evaluation [25]: Miss Rate (MR) over False Positive Per Image (also denoted as $\text{MR}^{\text{-2}}$). It calculates the average miss rate under the 9 FPPI values which are sampled uniformly in the logarithmic interval. The lower values indicate better performance.
IV-B2 Average Precision (AP)
@@
For FLIR and LLVIP datasets, we employ the commonly used object detection metric Average Precision (AP). The positive and negative samples should be divided according to the correctness of classification and Intersection over the Union (IoU) threshold. The mean Average Precision (mAP) represents the average of AP under all categories. The $\text{mAP}_{50}$ metric represents the mean AP under IoU=0.50 and The mAP metric represents the mean AP under IoU ranges from 0.50 to 0.95 with a stride of 0.05. The higher value of the mAP metrics means better performance.
@@
TABLE II: Camparison the performance (mAP, in%) on the FLIR and LLVIP dataset. The best results are highlighted in red and the second-place are highlighted in blue.
@@
      SSD [53]       VGG16       65.5       29.6       90.2       53.5
      RetinaNet [54]       ResNet50       66.1       31.5       94.8       55.1
      Cascade R-CNN [55]       ResNet50       71.0       34.7       95.0       56.8
      Faster R-CNN [47]       ResNet50       74.4       37.6       94.6       54.5
      DDQ-DETR [56]       ResNet50       73.9       37.1       93.9       58.6       IR
      SSD [53]       VGG16       52.2       21.8       82.6       39.8
      RetinaNet [54]       ResNet50       51.2       21.9       88.0       42.8
      Cascade R-CNN [55]       ResNet50       56.0       24.7       88.3       47.0
      Faster R-CNN [47]       ResNet50       64.9       28.9       87.0       45.1
      DDQ-DETR [56]       ResNet50       64.9       30.9       86.1       46.7       RGB
      Halfway fusion [17]       VGG16       71.5       35.8       91.4       55.1
      GAFF [57]       ResNet18       74.6       37.4       94.0       55.8
      ProbEn [58]       ResNet50       75.5       37.9       93.4       51.5
      LGADet [59]       ResNet50       74.5       -       -       -
      CSAA [60]       ResNet50       79.2       41.3       94.3       59.2
@@
IV-C Implementation Details
All the experiments are implemented in the mmdetection [61] toolbox and conducted on a single GPU NVIDIA GeForce RTX 3090. We use ResNet-50 as the backbone. During training, the batch size is set to 4 and the SGD optimizer is employed with a momentum of 0.9 and the weight decay of $1\times 10^{-4}$. The input image resolution is $640\times 512$ for FLIR and KAIST and $1280\times 1024$ for LLVIP. All experiments are trained for 12 epochs with an initial learning rate of $1\times 10^{-2}$. We use random flipping with a probability of 0.5 to increase input diversity.
IV-D Comparison with State-of-the-Art Methods
IV-D1 Comparision on the KAIST Dataset
The experiments are conducted on the KAIST under ‘All-dataset’ settings [25, 18]. We compare the proposed RSDet with twelve SOTA multispectral object detection methods on the KAIST dataset.As shown in Table I, We calculate the $\text{MR}^{\text{-2}}$ at IoU values of 0.5 and 0.7, respectively.
When IoU=0.5, according to Table I(a), the RSDet performs best on the ‘All’, ‘Day’, and ‘Night’ conditions and the other four of six subsets (‘Near’, ‘Medium’, ‘Far’, ‘None’). And rank second in two subsets (Partial and Heavy). Specifically, the RSDet takes the lead in the subsets of different scales (‘None’, ‘Medium’, ‘Far’), especially in the ‘Far’ subsets, the RSDet surpasses the second-best by about 10%, which is attributed to the mixture of scale-aware experts in the DFS module, Enhancing perceptual ability for targets at different distances. Besides, on the whole test dataset (‘All’), the RSDet surpasses the second-best by about 3%, which indicates the satisfactory accuracy of RSDet. The Log-average Miss Rate over the False Positive Per Image (FPPI) curve of all detectors under IoU=0.5 is shown in Figure 5(a).
When IoU=0.7, according to Table I(b), the RSDet performs best on all the different settings. As IoU increases, the requirement for detection accuracy becomes higher. In this scenario, the superiority of our method becomes even more pronounced. For instance, when IoU is 0.5, all the recent years’ methods MR is 0 under the ’Near’ subset. However, when IoU increases to 0.7, the miss rate of other methods increases more significantly than RSDet, indicating that our proposed method detection results more accurately. The Log-average Miss Rate over the False Positive Per Image (FPPI) curve of all detectors under IoU=0.7 is also shown in Figure 5(b). The above experiment results on the KAIST dataset prove the superiority of our method.
TABLE III: Ablation study on each module result (mAP, in%) on the FLIR and LLVIP dataset. The best results are highlighted in bold.
@@ IV-D1 Comparision on the KAIST Dataset
IV-D2 Comparision on the FLIR and LLVIP Dataset
The quantitative results of the different methods on the FLIR and LLVIP datasets are shown in Table II. We compare the proposed RSDet with five SOTA multispectral object detection methods and five single-modality object detection methods. From Table II, it can be observed that the experimental results of fusion methods are generally better than those of single-modal methods in the FLIR dataset. However, due to the limited light conditions, the RGB features interfere with the IR features in the LLVIP dataset, resulting in the detection results of the multispectral detector being inferior to the IR modality. However, our method effectively addresses this issue with the well-designed coarse-to-fine fusion strategy and achieves state-of-the-art performance on the both FLIR and LLVIP datasets.
Figure 7: Qualitative comparison of RSDet with other SOTA methods on the Kaist datasets. From left to right are the Ground Truth on IR and RGB image, MBNet [20], TSFADet [14], CAGTDet [31] and Ours.
Figure 8: Visualization of detection results of our method. The first two rows are from the FLIR dataset and the last row is from the LLVIP dataset. From left to right are the infrared image, the Ground Truth bounding box, and our detection results.
IV-E Ablation Study
IV-E1 Ablation on each component
To verify the effectiveness of the RSR and DFS modules, we ablate each module from RSDet. As shown in Table III, only by adding the RSR module does the performance increase by 1.2% mAP on the FLIR dataset and 1.6% mAP on the LLVIP dataset respectively, and by only adding the DFS module, the metrics show a significant increase on the FLIR and LLVIP dataset, the performance increase by 3.6% mAP on the FLIR dataset and 4.0% mAP on the LLVIP dataset, respectively. Further adding both the RSR and DFS modules, the metrics achieve the best level. The above results indicate the effectiveness of these modules.
IV-E2 Ablation on different K
We further study the impact of different values of $K$ in the RSR module. We conduct the experiments by modifying the value of $K$ ranging from 100 to 400 on the FLIR and LLVIP datasets. As shown in Figure 6, our detector achieves the highest detection performance when $K=320$. If the value of $K$ is increased or decreased, the performance of the detector will decrease, which also proves the negative impact of the redundant spectrum for the detector.
Figure 9: Example of intermediate results of RSR module on FLIR (1,2 rows) and LLVIP (3,4 rows) datasets. “Before” represents the input images, “After” represents the filtered image after the RSR module, and “SNR” is the signal-to-noise ratio. We further visualize the learned filter and the filtered image content in the second column and third column respectively. The last column is the comparison of the background and the foreground region.
Figure 10: tSNE visualization [62] of the modality-specific and shared features. (a) ‘w/o RSR’ represents without RSR module. (b) ‘w RSR’ represents with RSR module.
IV-F Visualization Results
IV-F1 Kaist dataset
We also visualize the object detection results of our RSDet with other state-of-the-art methods on the Kaist testing dataset. From the first two rows of Figure 7, a common issue observed across the MBNet [20], TSFADet [14] and CAGTDet [31] methods is the missed detection of small objects. These results underscore the effectiveness of our DFS module, which can dynamically select different scale features between two modalities. Besides, by cooperating with the RSR module, our method can extract object features more effectively under complex lighting conditions and improve the performance of the detector, which is reflected in the confidence of the bounding box in the last row of Figure 7. The above results verify the effectiveness of our method.
IV-F2 FLIR and LLVIP dataset
In Figure 8, we visualize the object detection results of our method on the FLIR and LLVIP testing dataset. From Figure 8, we can see that our detection results include almost all bounding boxes of the Ground Truth labels with a high degree of consistency. Furthermore, our method can also detect some small objects that are miss-labeled in the Ground Truth, which demonstrates the superiority of our method in multi-scale object detection.
@@ IV-F2 FLIR and LLVIP dataset
IV-F3 RSR module
We visualize the intermediate results of the RSR module as shown in Figure 9. The second column visualizes the filter used in the frequency domain. Intuitively, both high- and low-frequency information will be filtered. We further transform the filtered information into the spatial domain, visualized in the third column. We can find that the removed information is mainly concentrated in the background region, which illustrates that the RSR module filters out some background redundant noise irrelevant to the detection task. For further objective comparison, we use the signal-to-noise ratio (SNR) to measure the differences, which is calculated using the variance method: $SNR=20\times\log(\frac{\mu}{\sigma^{2}})$. It can be observed that SNR increases after using the RSR module, indicating the effectiveness of the RSR module.
We also conduct tSNE visualizations [62] of shared $C_{\text{sha}}$ and specific features ($C_{\text{I-spe}}$ and $C_{\text{V-spe}}$) on FLIR dataset. As shown in Figure 10, we have observed that without the RSR module, there is still a considerable mix of feature points in the shared and specific features, leading to difficulty in selecting the desired specific features in DFS. After adding the RSR module, we have noted a significant decrease in the number of mixed features, which verifies that removing redundant spectra is beneficial to feature disentanglement and, thus more effective in obtaining fused features in the DFS module.
IV-F4 DFS module
@@
V Conclusion
In this paper, we presented a new coarse-to-fine perspective to fuse visible and infrared modality features. Specifically, a Redundant Spectrum Removal (RSR) module is first designed to coarsely filter out the irrelevant spectrum, and then a Dynamic Feature Selection (DFS) module is proposed to finely select the desired features for the RGB-IR final feature fusion process. we constructed a new object detector called Removal and Selection Detector (RSDet) to evaluate its effectiveness and versatility. Extensive experiments on three public RGB-IR detection datasets demonstrated that our method can effectively facilitate complementary fusion and achieve state-of-the-art performance. We believe that our method can be applied to various studies in the RGB-IR feature fusion tasks.